# Quantization Error Propagation: Revisiting Layer-Wise Post-Training Quantization

### v2

@@ A.4 Relationship of QEP Correction and Ridge Regularization
We formally establish a rigorous mathematical connection between the Quantization Error Propagation (QEP) correction parameter $\alpha_{l}$ and the ridge regularization parameter $\lambda$. Specifically, we show that tuning the QEP parameter $\alpha_{l}$ can be equivalently interpreted as adjusting the strength of ridge regularization with parameter $\lambda$. Furthermore, we prove the monotone inverse relationship between these two parameters.
Proposition A.6.
@@
$\widehat{{\bm{W}}}_{l}^{\ast}(\alpha_{l})={\bm{W}}_{l}({\bm{I}}+\alpha{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1})$ (74)
whereas the unique minimiser of the ridge objective
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{R}}^{n_{l}\times m}}f(\widehat{{\bm{W}}}_{l}),~{}~{}f(\widehat{{\bm{W}}}_{l})=\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}^{2}+\lambda_{l}\|{\bm{W}}_{l}-\widehat{{\bm{W}}}_{l}\|_{F},~{}~{}\lambda_{l}\geq 0,$ (75)
@@
$\widehat{{\bm{W}}}^{\ast}_{l}(\lambda_{l})=\widehat{{\bm{W}}}_{l}\left({\bm{I}}+{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}})^{-1}\right).$ (76)
@@
Thus, decreasing $\lambda$ from +$\infty$+ ∞ to 00 corresponds to increasing $\alpha_{l}$ from 00 to $1$.
@@
$\widehat{{\bm{W}}}_{l}(\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}})={\bm{W}}_{l}({\bm{X}}_{l}\widehat{{\bm{X}}}_{l}+\lambda{\bm{I}}),$ (82)
@@
Diagonalise $\widehat{{\bm{H}}}_{l}={\bm{U}}{\bm{\Gamma}}{\bm{U}}^{\top}$ with $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{d_{l}})$. Then ${\bm{G}}(\alpha_{l})={\bm{U}}\alpha{\bm{\Gamma}}^{-1}{\bm{U}}^{\top}$ has eigenvalues $\nicefrac{{\alpha_{l}}}{{d}}$, which increase strictly with $\alpha_{l}$, while ${\bm{R}}(\lambda_{l})={\bm{U}}({\bm{\Gamma}}+\lambda{\bm{I}})^{-1}{\bm{U}}^{\top}$ has eigenvalues $\nicefrac{{1}}{{\gamma_{i}+\lambda}}$, which decrease with $\lambda_{l}$, which means Loewner relations follow directly.
@@
which means $\alpha^{\prime}(\lambda)<0,\forall\lambda\geq 0$. Thus, $\alpha(\cdot)$ is strictly decreasing on $[0,\infty)$. One has
@@
$\alpha(\cdot)$ is strictly decreasing from $1$ to 00 and smooth on $[0,\infty)$. Because $\alpha$ is continous, strictly decreasing, it is a bijection from $[0,+\infty)$ onto $(0,1]$. By construction
@@
Thus, decreasing $\lambda$ from +$\infty$+ ∞ to 00 corresponds to increasing $\alpha_{l}$ from 00 to $1$.
A.5 Technical Lemma
Lemma A.7.
Let ${\bm{Z}}\in{\mathbb{R}}^{m\times n}$ be arbitrary and let $P\in{\mathbb{R}}^{n\times n}$ be an orthogonal projection, i.e. ${\bm{P}}^{2}={\bm{P}}$ and ${\bm{P}}^{\top}={\bm{P}}$. For every pair $0\leq\alpha^{\prime}\leq\alpha\leq 1$,
@@ A.5 Technical Lemma
$f(\alpha)=\mathrm{Tr}\left[({\bm{I}}-\alpha{\bm{P}}){\bm{Z}}^{\top}{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\right]=\|{\bm{Z}}\|_{F}^{2}-\alpha(2-\alpha)\underbrace{\mathrm{Tr}({\bm{Z}}^{\top}{\bm{Z}}{\bm{P}})}_{t\geq 0}.$ (88)
Hence $f^{\prime}(\alpha)=-(2-\alpha)t\leq 0$ on $[0,1]$, so $f(\alpha)$ is non-increasing. Taking square roots yields the first inequality in (87). Setting $\alpha^{\prime}=0$ gives the second inequality: $\|{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\|_{F}\leq\|{\bm{Z}}\|_{F}$. ∎
Lemma A.8.
Fix a layer index $l\in\{1,\dots,L\}$ and let ${\bm{P}}_{l}\in\mathbb{R}^{m\times m}$ be an orthogonal projection, ${\bm{P}}_{l}^{2}={\bm{P}}_{l}={\bm{P}}_{l}^{\top}$. For a given matrix ${\bm{Z}}_{l}\in\mathbb{R}^{d_{l}\times m}$ define the pre-activation that uses a parameter $\alpha\in[0,1]$ by
$\widehat{{\bm{Y}}}_{l}^{\alpha}\coloneqq{\bm{Z}}_{l}\left({\bm{I}}-\alpha{\bm{P}}_{l}\right)\in\mathbb{R}^{d_{l}\times m},~{}\widehat{{\bm{X}}}_{l}^{\alpha}\coloneqq\sigma_{l}\left(\widehat{{\bm{Y}}}_{l}^{\alpha}\right).$ (89)
Assume the non-linearity $\sigma_{l}$ is $1$-Lipschitz with respect to the Frobenius norm,
$\|\sigma_{l}({\bm{A}})-\sigma_{l}({\bm{B}})\|_{F}\leq\|{\bm{A}}-{\bm{B}}\|_{F},~{}~{}\forall{\bm{A}},{\bm{B}}.$ (90)
Then, for any $0\leq\alpha^{\prime}\leq\alpha\leq 1$,
$\left\|\widehat{{\bm{X}}}_{l}^{\alpha}\right\|_{F}\leq\left\|\widehat{{\bm{X}}}_{l}^{\alpha^{\prime}}\right\|_{F}.$ (91)
Proof.
Lemma A.8 states that for every fixed matrix ${\bm{Z}}$ the map $\alpha\mapsto\|{\bm{Z}}({\bm{I}}-\alpha{\bm{P}}_{l})\|_{F}$ is non-increasing on $[0,1]$. Applying it to ${\bm{Z}}={\bm{Z}}_{l}$ yields
$\left\|\widehat{{\bm{Y}}}_{l}^{\alpha}\right\|_{F}=\left\|{\bm{Z}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})\right\|_{F}\leq\left\|{\bm{Z}}_{l}({\bm{I}}-\alpha^{\prime}{\bm{P}}_{l})\right\|_{F}=\left\|\widehat{{\bm{Y}}}_{l}^{\alpha^{\prime}}\right\|_{F}.$ (92)
Because $\sigma_{l}$ is $1$-Lipschitz, it cannot increase the Frobenius norm of any single argument: $\|\sigma_{l}({\bm{A}})\|_{F}\leq\|{\bm{A}}\|_{F}$. Hence
$\bigl{\|}\widehat{{\bm{X}}}_{l}^{\alpha}\bigr{\|}_{F}=\bigl{\|}\sigma_{l}(\widehat{{\bm{Y}}}_{l}^{\alpha})\bigr{\|}_{F}\leq\bigl{\|}\widehat{{\bm{Y}}}_{l}^{\alpha}\bigr{\|}_{F}\leq\bigl{\|}\widehat{{\bm{Y}}}_{l}^{\alpha^{\prime}}\bigr{\|}_{F}\leq\bigl{\|}\sigma_{l}(\widehat{{\bm{Y}}}_{l}^{\alpha^{\prime}})\bigr{\|}_{F}=\bigl{\|}\widehat{{\bm{X}}}_{l}^{\alpha^{\prime}}\bigr{\|}_{F},$ (93)
which proves the claim. ∎
Appendix B Additional Implementation Details
B.1 Damping for Hessian
A standard numerical issue in PTQ arises when the Hessian matrix $\widehat{{\bm{H}}}_{l}$ is ill-conditioned or even singular, making its inversion unstable or undefined. Following GPTQ [Frantar et al., 2022], we address this issue using a damping strategy that adds a small scalar value $\lambda$ to the diagonal elements of $\widehat{{\bm{H}}}_{l}$ to ensure positive definiteness. In our implementation, we set $\lambda$ to the mean of the diagonal elements of $\widehat{{\bm{H}}}_{l}$, offering a simple yet effective way to stabilize the inversion process.
@@ B.1 Damping for Hessian
Appendix C Additional Experiments
C.1 Additional Perplexity Results
Due to space constraints, the main text reports perplexity results only for the WikiText-2 dataset. Here, we provide additional results for PTB (Table 6) and C4 (Table 7), alongside extended WikiText-2 results (Table 5). These supplementary results further validate that QEP consistently improves PTQ performance, especially in low-bit quantization scenarios.
@@ C.1 Additional Perplexity Results
C.2 Detailed Accuracy Results for Individual Tasks
@@
C.3 Stability of QuIP Results Across Random Seeds
We assess the stability of QuIP-only experiments by averaging five independent runs per configuration. Model sizes, quantization levels, and benchmarks match the main Experiments section. Figure 3 plots QuIP with or without QEP at three quantization levels. Each marker is the mean of five seeds, and the error bars show the standard error of the mean. The top row gives perplexity on WikiText 2; the bottom row reports mean normalized accuracy on ARC easy, PIQA, and StoryCloze. Seed-to-seed variation is small and does not change the main conclusions. QEP-QuIP keeps its advantage, especially at INT3 and INT2. The main text lists the best seed per configuration for consistency with past work. This appendix confirms that the gains are not seed-specific but robust and reproducible, supporting using QEP.

### v3

@@ B.4 Relationship of QEP Correction and Ridge Regularization
We formally establish a rigorous mathematical connection between the Quantization Error Propagation (QEP) correction parameter $\alpha_{l}$ and the ridge regularization parameter $\lambda$. We show that tuning the QEP parameter $\alpha_{l}$ is equivalent to adjusting the strength of ridge regularization with parameter $\lambda$. We prove the monotone inverse relationship between these two parameters.
Proposition B.7.
@@
$\widehat{{\bm{W}}}_{l}^{\ast}(\alpha_{l})={\bm{W}}_{l}({\bm{I}}+\alpha{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1}).$ (100)
The unique minimizer of the ridge objective
$\min_{\widehat{{\bm{W}}}_{l}\in{\mathbb{R}}^{n_{l}\times d_{l}}}f(\widehat{{\bm{W}}}_{l}),~~f(\widehat{{\bm{W}}}_{l})=\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}^{2}+\lambda_{l}\|{\bm{W}}_{l}-\widehat{{\bm{W}}}_{l}\|_{F}^{2},~~\lambda_{l}\geq 0,$ (101)
@@
$\widehat{{\bm{W}}}^{\ast}_{l}(\lambda_{l})={\bm{W}}_{l}\left({\bm{I}}+{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}})^{-1}\right).$ (102)
@@
Thus, decreasing $\lambda$ from $\infty$ to 0 corresponds to increasing $\alpha_{l}$ from 0 to $1$.
@@
$\widehat{{\bm{W}}}_{l}(\widehat{{\bm{H}}}_{l}+\lambda{\bm{I}})={\bm{W}}_{l}({\bm{X}}_{l}\widehat{{\bm{X}}}_{l}^{\top}+\lambda{\bm{I}}),$ (108)
@@
Diagonalise $\widehat{{\bm{H}}}_{l}={\bm{U}}{\bm{\Gamma}}{\bm{U}}^{\top}$ with $\Gamma=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{d_{l}})$. Then ${\bm{G}}(\alpha_{l})={\bm{U}}\alpha_{l}{\bm{\Gamma}}^{-1}{\bm{U}}^{\top}$ has eigenvalues $\nicefrac{{\alpha_{l}}}{{\gamma_{l}}}$, which increase strictly with $\alpha_{l}$, while ${\bm{R}}(\lambda_{l})={\bm{U}}({\bm{\Gamma}}+\lambda{\bm{I}})^{-1}{\bm{U}}^{\top}$ has eigenvalues $\nicefrac{{1}}{{\gamma_{i}+\lambda}}$, which decrease with $\lambda_{l}$, which means Loewner relations follow directly.
@@
which means $\alpha^{\prime}(\lambda)<0,\forall\lambda\geq 0$. Thus, $\alpha(\cdot)$ is strictly decreasing on $0,\infty)$. One has
@@
$\alpha(\cdot)$ is strictly decreasing from $1$ to 0 and smooth on $0,\infty)$. Because $\alpha$ is continous, strictly decreasing, it is a bijection from $0,+\infty)$ onto $(0,1]$. By construction
@@
Thus, decreasing $\lambda$ from $\infty$ to 0 corresponds to increasing $\alpha_{l}$ from 0 to $1$.
B.5 Technical Lemma
Lemma B.8.
Let ${\bm{Z}}\in{\mathbb{R}}^{m\times n}$ be arbitrary, and let $P\in{\mathbb{R}}^{n\times n}$ be an orthogonal projection, i.e., ${\bm{P}}^{2}={\bm{P}}$ and ${\bm{P}}^{\top}={\bm{P}}$. For every pair $0\leq\alpha^{\prime}\leq\alpha\leq 1$,
@@ B.5 Technical Lemma
$f(\alpha)=\mathrm{Tr}\left[({\bm{I}}-\alpha{\bm{P}}){\bm{Z}}^{\top}{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\right]=\|{\bm{Z}}\|_{F}^{2}-2\alpha(1-\alpha)\underbrace{\mathrm{Tr}({\bm{Z}}^{\top}{\bm{Z}}{\bm{P}})}_{t\geq 0}.$ (114)
Thus, $f^{\prime}(\alpha)=-(2-\alpha)t\leq 0$ on $[0,1]$ indicates that $f(\alpha)$ is non-increasing. Taking square roots yields the first inequality in Eq. (113). Setting $\alpha^{\prime}=0$ yields the second inequality: $\|{\bm{Z}}({\bm{I}}-\alpha{\bm{P}})\|_{F}\leq\|{\bm{Z}}\|_{F}$. ∎
Appendix C Additional Implementation Details
C.1 Damping for Hessian
A standard numerical issue in PTQ arises when the Hessian matrix $\widehat{{\bm{H}}}_{l}$ is ill-conditioned or singular, rendering its inversion unstable or undefined. Following GPTQ [frantar2022gptq], we resolve this issue by employing a damping strategy that adds a small scalar value $\lambda$ to the diagonal elements of $\widehat{{\bm{H}}}_{l}$ to ensure positive definiteness. In our implementation, we set $\lambda$ to the mean of the diagonal elements of $\widehat{{\bm{H}}}_{l}$, providing a straightforward yet effective method to stabilize the inversion process.
@@ C.1 Damping for Hessian
Appendix D Additional Experiments
D.1 Additional Perplexity Results
Due to space constraints, the main text reports perplexity results solely for the WikiText-2 dataset. Here, we provide additional results for PTB (Table 6) and C4 (Table 7), along with extended results for WikiText-2 (Table 5). These supplementary results further validate that QEP consistently enhances PTQ performance, particularly in low-bit quantization scenarios.
@@ D.1 Additional Perplexity Results
D.2 Detailed Accuracy Results for Individual Tasks
@@
D.3 Stability of QuIP Results Across Random Seeds
We assess the stability of QuIP-only experiments by averaging five independent runs per configuration. Model sizes, quantization levels, and benchmarks align with the main Experiments section. Figure 3 plots QuIP with or without QEP at three quantization levels. Each marker is the mean of five seeds, and the error bars show the standard error of the mean. The top row gives perplexity on WikiText 2; the bottom row reports mean normalized accuracy on ARC easy, PIQA, and StoryCloze. Seed-to-seed variation is small and does not change the main conclusions. QEP-QuIP keeps its advantage, especially at INT3 and INT2. The main text lists the best seed per configuration for consistency with past work. This appendix confirms that the gains are not seed-specific but robust and reproducible, supporting using QEP.
Table 11: WikiText-2 perplexity for LLaMA-2-7B at different bit-widths. NaN denotes divergence.
Method INT4 INT3 INT2
RTN+QEP 6.017 17.309 97153.266
GPTQ+QEP 5.933 7.898 7214.328
AWQ+QEP 5.756 11.131 229888.406
QuIP+QEP 5.753 6.154 11.972
OmniQuant 5.880 7.065 NaN
D.4 Comparison with OmniQuant Baseline
For completeness, we compare QEP-enhanced layer-wise PTQ with block-wise OmniQuant [shao2023omniquant] on LLaMA-2-7B using WikiText-2 perplexity; lower values indicate better performance. As shown in Table 11, QuIP+QEP achieves the lowest perplexity at INT4/INT3 and remains stable at INT2, while OmniQuant diverges. These findings align with recent PTQ benchmarks indicating OmniQuant’s underperformance relative to layer-wise PTQ [zhao2025benchmarking].
[…]